the length of a rectangle is one and a half times the width. the perimeter is twenty five centimetres. find the length and width
Answer by TimothyLamb(4379)

L = 1.5w
2L + 2w = 25
---
2L + 2w = 25
2(1.5w) + 2w = 25
3w + 2w = 25
5w = 25
w = 5
---
L = 1.5*5
L = 7.5
---
answer:
L = 7.5 cm
w = 5 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php